the perimeter of a rectangular yard is 720. if the length is 25 feet greater than its width what are the dimensions of the yard?
Answer by ewatrrr(24785)


 
Hi,

P = 2L + 2w = 720 0r L + w = 360
Question States*** L = w + 25
(w+25) + w = 360
2w = 335
w= 167.5, the width. the length is 192.5
CHECKING our answer***
167.5 + 192.5 = 360
Wish You the Best in your Studies.